Given that log_2 x=5, evaluate:
log_2 (x/2)

Answer by stanbon(75887)

Given that log_2 x=5, evaluate:
-----
log_2(4x^2) = log2 (4) + 2log2 (x)
= 2 + 2*5
= 12
================
Cheers,
Stan H.